OpenStudy (anonymous):

Let f=(x) = x2 – 16 and g(x) = x + 4. Find f/g and its domain.

OpenStudy (binary3i):

factorize f(x) so f(x)=(x-4)(x+4) now f(x)/g(x)=(x-4)(x+4)/(x+4)=(x-4) its domain is all real no. in your options you have mentioned that x not equal to ither 4 or -4 but the domain is all the real no. with 4 and -4

OpenStudy (anonymous):

OK thx for the help :p